# Kernelized Concept Erasure

### v5

@@ 3.3.1 Convexifying the Objective
$\min_{{\boldsymbol{\theta}}\in\Theta}\,\max_{P\in\mathcal{F}_{k}}\,\sum_{n=1}^{N}\ell\Big{(}y_{n},\,\langle{\boldsymbol{\theta}},P\boldsymbol{\widetilde{\Phi}}(\mathbf{x}_{n})\rangle\Big{)}$ (10)
which is identical to the linear game in Eq. 3, except that it uses the transformed features $\boldsymbol{\widetilde{\Phi}}(\mathbf{x}_{n})$ to approximate the true feature map $\boldsymbol{\Phi}$.
While neither the game in Eq. 3 nor the version of Eq. 3 with the Nyström approximation given in Eq. 10 is convex–concave, we are still able to approximately solve the game in practice. The runtime of a gradient-based optimization procedure for this game over $T$ epochs is now $\mathcal{O}\left(TN^{2}+N^{3}\right)$. [^5] Calculating the Nyström features entails computing the eigendecomposition of the Gram matrix, which is $\mathcal{O}\left(N^{3}\right)$. Then, we calculate the objective over all $N$ examples for $T$ epochs, which is $\mathcal{O}\left(TN^{2}\right)$. While this is an improvement over $\mathcal{O}\left(TN^{4}\right)$, it is still not fast enough to be of practical use.
@@ 3.3.2 Improving the Runtime
The runtime bottleneck of the game in Eq. 10 is the $\mathcal{O}\left(N^{3}\right)$ time it takes to compute the eigendecomposition of the Gram matrix $K$. Under the assumption that $\mathrm{rank}\left(K\right)=L$, we can improve this bound to $\mathcal{O}\left(L^{3}+L^{2}N\right)$ (Drineas and Mahoney, 2005). For the case that $L\ll N$, this is a substantial improvement. Moreover, it implies that the approximation feature map $\widetilde{\boldsymbol{\Phi}}(\mathbf{x}_{n})\in\mathbb{R}^{L}$. After $T$ steps of optimization, the runtime is now $\mathcal{O}\left(TL^{2}+L^{3}+L^{2}N\right)$, which is fast enough to be useful in practice. A natural question to ask is what happens if we apply the Nyström approximation, thereby assuming $\mathrm{rank}\left(K\right)=L$, but in practice $\mathrm{rank}\left(K\right)>L$? In this case, we are effectively computing a low-rank approximation of the kernel matrix. Several bounds on the accuracy of this approximation have been proven in the literature (Drineas and Mahoney, 2005; Jin et al., 2013; Nemtsov et al., 2016); we refer the reader to these works for more details on the approximation error.
@@ 3.4 Pre-image Mapping
After solving the game in Eq. 10, we obtain a projection matrix $P$ that neutralizes the concept inside an (approximated) RKHS. In other words, we have a function $r(\cdot)$ that prevents a classifier from predicting a concept label from the representation $\widetilde{\boldsymbol{\Phi}}(\mathbf{x}_{n})$. However, for many applications, we want a version of the input $\mathbf{x}_{n}$ in the original space with the concept neutralized, i.e., we want $r(\mathbf{x}_{n})$. Neutralization in the original space requires solving the pre-image problem (Mika et al., 1998). In the case of Nyström features, we seek a mapping $P\widetilde{\boldsymbol{\Phi}}(\mathbf{x})\mapsto\mathbf{x}$, which is a mapping from $\mathbb{R}^{L}$ to $\mathbb{R}^{D}$, that projects the neutralized features back into the input space.
In practice, this task can also be performed via a mapping $\mathbf{x}\mapsto\mathbf{x}$ from $\mathbb{R}^{D}\mapsto\mathbb{R}^{D}$, which learns to reproduce in the input space the transformation that $P$ performs in the RKHS. We choose the latter approach, and train a multilayer perceptron (MLP) $f_{\lambda}(\cdot):\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$. [^6] This mapping may not always be a function because there may be more than one point in the RKHS that corresponds to a point in the original space. To estimate the parameters $\lambda$ of the MLP $f_{\lambda}(\cdot)$, we optimize a two-termed objective over all points $\mathbf{x}_{n}$:
$\operatorname*{\mathop{\text{argmin}}}_{\lambda\in\Lambda}$ ${\left|\left|P\boldsymbol{\widetilde{\Phi}}(\mathbf{x}_{n})-\boldsymbol{\boldsymbol{\widetilde{\Phi}}}(f_{\lambda}(\mathbf{x}_{n}))\right|\right|_{2}^{2}}$ (11)
$\quad+\left|\left|(I-P)\boldsymbol{\widetilde{\Phi}}(f_{\lambda}(\mathbf{x}_{n}))\right|\right|_{2}^{2}$
where $\Lambda$ is the parameter space. The first term encourages $f_{\lambda}(\cdot)$ to perform the same transformation in the input space as $P$ does in the RKHS. The second term ensures that $P$ has no effect on the RKHS features computed on the neutralized $f_{\lambda}(\mathbf{x}_{n})$.
@@ 4 Experimental Setup
  • Poly : $\kappa(\mathbf{x},\mathbf{y})={(\gamma\mathbf{x}^{\top}\mathbf{y}+\alpha)}^{\texttt{d}}$
  • RBF : $\kappa(\mathbf{x},\mathbf{y})=\exp\left(-\gamma||\mathbf{x}-\mathbf{y}||_{2}^{2}\right)$
  • Laplace : $\kappa(\mathbf{x},\mathbf{y})=\exp\left(-\gamma||\mathbf{x}-\mathbf{y}||_{1}\right)$
  • Linear : $\kappa(\mathbf{x},\mathbf{y})=\mathbf{x}^{\top}\mathbf{y}$
  • Sigmoid : $\kappa(\mathbf{x},\mathbf{y})=\mathrm{tanh}(\gamma\mathbf{x}^{\top}\mathbf{y}+\alpha)$
@@
Solving the minimax game.
We solve the relaxed adversarial game given in Eq. 10 by alternate gradient-based optimization (Goodfellow et al., 2016). Concretely, we alternate between updating the predictor’s parameters ${\boldsymbol{\theta}}$ and the projection matrix $P$. Updates to ${\boldsymbol{\theta}}$ are performed with gradient descent, and updates to $P$ are performed with gradient ascent, including a projection onto the Fantope to ensure that the constraint is met. For the Fantope projection step, we use Vu et al.’s (2013) algorithm, the details of which are restated in Ravfogel et al. (2022).
@@ 5 Effect on Concept Encoding
5.1 Pre-image Gender Recovery: Same Adversary
5.2 Pre-image Gender Recovery: Transfer Between Kernels
@@ Results.
In Table 3, we report the test statistic and the $p$ value for the third test using the names of scientific and artistic fields to represent gender-biased words. [^10] See § A.6 for results of the other two biased word association tests. For all kernels, we observe a significant drop in the test statistic, from the original value of 1.56 to around 0.74. This suggests that the intervention significantly decreases the association between female and male names and stereotypically biased words. Notably, the reduction is similar for all kernels, including the linear one. While non-linear erasure is more effective in neutralizing gender against adversarial recovery, linear and non-linear methods perform equally well according to this bias association test. This finding highlights the importance of measuring different manifestations of a concept when using concept neutralization as a bias mitigation method.
@@ Appendix A Related Work
A.1 A Representer Lemma for Kernelized Minimax Games
@@
$\max_{\boldsymbol{w},\in\mathcal{H}}\min_{{\boldsymbol{\theta}}\in\mathcal{H}}\sum_{n=1}^{N}\ell\left(y_{n},\,\langle{\boldsymbol{\theta}},\left(\mathrm{I}-\frac{{\boldsymbol{w}\boldsymbol{w}^{\top}}}{\boldsymbol{w}^{\top}\boldsymbol{w}}\right)\boldsymbol{\Phi}(\mathbf{x}_{n})\rangle\right)$ (12)
@@
$=\boldsymbol{\Phi}(\mathbf{x}_{n})-\frac{\boldsymbol{w}_{U}\boldsymbol{w}_{U}^{\top}}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}\boldsymbol{\Phi}(\mathbf{x}_{n})-\frac{\boldsymbol{w}_{\perp U}\boldsymbol{w}_{\perp U}^{\top}}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}\boldsymbol{\Phi}(\mathbf{x}_{n})$ (13d)
$=\boldsymbol{\Phi}(\mathbf{x}_{n})-\boldsymbol{w}_{U}\frac{\boldsymbol{w}_{U}^{\top}\boldsymbol{\Phi}(\mathbf{x}_{n})}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}$ (13e)
@@
$\Big{\langle}{\boldsymbol{\theta}},\left(\mathrm{I}-{\boldsymbol{w}\boldsymbol{w}^{\top}}\right)\boldsymbol{\Phi}(\mathbf{x}_{n})\Big{\rangle}={\boldsymbol{\theta}}_{U}^{\top}\boldsymbol{\Phi}(\mathbf{x}_{n})-{\boldsymbol{\theta}}_{U}^{\top}\boldsymbol{w}_{U}\frac{\boldsymbol{w}_{U}^{\top}\boldsymbol{\Phi}(\mathbf{x}_{n})}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}$ (14)
Thus, for any ${\boldsymbol{\theta}},\boldsymbol{w},\in\mathcal{H}$ there exists a ${\boldsymbol{\theta}}_{U},\boldsymbol{w}_{U}\in U$ that yields the same value of the objective as ${\boldsymbol{\theta}},\boldsymbol{w}$. Now, we can parameterize ${\boldsymbol{\theta}}_{U}$ and $\boldsymbol{w}_{U}$ as
@@
${\boldsymbol{\theta}}_{U}$ $=\sum_{n=1}^{N}\beta_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})$ (16)
for real coefficients $\boldsymbol{\alpha}\in\mathbb{R}^{N}$ and $\boldsymbol{\beta}\in\mathbb{R}^{N}$. We conclude that for any local optimum ${\boldsymbol{\theta}}^{*},\boldsymbol{w}^{*}\in\mathcal{H}$, the projection of ${\boldsymbol{\theta}}^{*}$ and $\boldsymbol{w}^{*}$ onto $U$ yields a local optimum with the same value. ∎
We note that under regularity conditions, i.e., certain compactness and convexity restrictions over the feasible sets and the loss function, the min and the max can be swapped as per the celebrated Von Neumann–Fan minimax theorem (Fan, 1953). For the aforementioned reasons, we believe Lemma 1 justifies the parameterizations used for ${\boldsymbol{\theta}},\boldsymbol{w}$, e.g., Eq. 19.
A.2 Kernelization of the Minimax Game
We show that the game Eq. 1 can be kernelized for the case $k=1$, i.e., a setting where the matrix $\mathcal{P}_{k}$ removes a one-dimensional subspace. Specifically, we will show that the product $\langle{\boldsymbol{\theta}},\mathrm{P}\boldsymbol{\Phi}(\mathbf{x}_{n})\rangle$ in Eq. 1 can be expressed as a function of the kernel $\kappa(\cdot,\cdot)$.

### v6

@@ 3.3.1 Relaxing the Objective
$\max_{\boldsymbol{w}\in\mathbb{R}^{D}}\,\min_{{\boldsymbol{\theta}}\in\Theta}\,\sum_{n=1}^{N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ell}\Big{(}y_{n},\,\langle{\boldsymbol{\theta}},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}_{\boldsymbol{w}}^{\bot}\boldsymbol{\widetilde{\Phi}}(\boldsymbol{x}_{n})\rangle\Big{)}$ (10)
which is identical to the linear game in Eq. 4, except that it uses the transformed features $\boldsymbol{\widetilde{\Phi}}(\boldsymbol{x}_{n})$ to approximate the true feature map $\boldsymbol{\Phi}$.
While neither the game in Eq. 4 nor the version of Eq. 4 with the Nyström approximation given in Eq. 10 is convex–concave, we are still able to approximately solve the game in practice. Specifically, we consider the convex relaxation of the set of rank-${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}K}$ orthogonal projection matrices given by Ravfogel et al. (2022). Then, after applying the relaxation, we apply a gradient-based optimization procedure. The runtime of this procedure for this game over $T$ epochs is now $\mathcal{O}\left(TN^{2}+N^{3}\right)$. [^5] Calculating the Nyström features entails computing the eigendecomposition of the Gram matrix, which is $\mathcal{O}\left(N^{3}\right)$. Then, we calculate the objective over all $N$ examples for $T$ epochs, which is $\mathcal{O}\left(TN^{2}\right)$. While this is an improvement over $\mathcal{O}\left(TN^{4}\right)$, it is still not fast enough to be of practical use.
@@ 3.3.2 Improving the Runtime
The runtime bottleneck of the game in Eq. 10 is the $\mathcal{O}\left(N^{3}\right)$ time it takes to compute the eigendecomposition of the Gram matrix $K$. Under the assumption that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{rank}}\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}}\right)=L$, we can improve this bound to $\mathcal{O}\left(L^{3}+L^{2}N\right)$ (Drineas and Mahoney, 2005). For the case that $L\ll N$, this is a substantial improvement. Moreover, it implies that the approximation feature map $\widetilde{\boldsymbol{\Phi}}(\boldsymbol{x}_{n})\in\mathbb{R}^{L}$. After $T$ steps of optimization, the runtime is now $\mathcal{O}\left(TL^{2}+L^{3}+L^{2}N\right)$, which is fast enough to be useful in practice. A natural question to ask is what happens if we apply the Nyström approximation, thereby assuming ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{rank}}\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}}\right)=L$, but in practice ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{rank}}\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}}\right)>L$? In this case, we are effectively computing a low-rank approximation of the kernel matrix. Several bounds on the accuracy of this approximation have been proven in the literature (Drineas and Mahoney, 2005; Jin et al., 2013; Nemtsov et al., 2016); we refer the reader to these works for more details on the approximation error.
@@ 3.4 Pre-image Mapping
After solving the game in Eq. 10, we obtain a projection matrix ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}$ that neutralizes the concept inside an (approximated) RKHS. In other words, we have a function $r(\cdot)$ that prevents a classifier from predicting a concept label from the representation $\widetilde{\boldsymbol{\Phi}}(\boldsymbol{x}_{n})$. However, for many applications, we want a version of the input $\boldsymbol{x}_{n}$ in the original space with the concept neutralized, i.e., we want $r(\boldsymbol{x}_{n})$. Neutralization in the original space requires solving the pre-image problem (Mika et al., 1998). In the case of Nyström features, we seek a mapping from feature space back to the original Euclidean space, i.e., a mapping from $\mathbb{R}^{L}$ to $\mathbb{R}^{D}$, that projects the neutralized features back into the input space.
In practice, this task can also be performed via a mapping from $\mathbb{R}^{D}\mapsto\mathbb{R}^{D}$, which learns to reproduce in the input space the transformation that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}$ performs in the RKHS. We choose the latter approach, and train a multilayer perceptron (MLP) $f_{\lambda}(\cdot)\colon\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$. [^6] This mapping may not always be a function because there may be more than one point in the RKHS that corresponds to a point in the original space. To estimate the parameters $\lambda$ of the MLP $f_{\lambda}(\cdot)$, we optimize a two-termed objective over all points $\boldsymbol{x}_{n}$:
$\operatorname*{\mathop{\text{argmin}}}_{\lambda\in\Lambda}$ ${\left|\left|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}\boldsymbol{\widetilde{\Phi}}(\boldsymbol{x}_{n})-\boldsymbol{\boldsymbol{\widetilde{\Phi}}}(f_{\lambda}(\boldsymbol{x}_{n}))\right|\right|_{2}^{2}}$ (11)
$\quad+\left|\left|({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{I}}-{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}})\boldsymbol{\widetilde{\Phi}}(f_{\lambda}(\boldsymbol{x}_{n}))\right|\right|_{2}^{2}$
where $\Lambda$ is the parameter space. The first term encourages $f_{\lambda}(\cdot)$ to perform the same transformation in the input space as ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}$ does in the RKHS. The second term ensures that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}$ has no effect on the RKHS features computed on the neutralized $f_{\lambda}(\boldsymbol{x}_{n})$.
@@ 4 Experimental Setup
  • Poly : ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x},\boldsymbol{y})={(\gamma\boldsymbol{x}^{\top}\boldsymbol{y}+\alpha)}^{\texttt{d}}$
  • RBF : ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x},\boldsymbol{y})=\exp\left(-\gamma||\boldsymbol{x}-\boldsymbol{y}||_{2}^{2}\right)$
  • Laplace : ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x},\boldsymbol{y})=\exp\left(-\gamma||\boldsymbol{x}-\boldsymbol{y}||_{1}\right)$
  • Linear : ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x},\boldsymbol{y})=\boldsymbol{x}^{\top}\boldsymbol{y}$
  • Sigmoid : ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x},\boldsymbol{y})=\mathrm{tanh}(\gamma\boldsymbol{x}^{\top}\boldsymbol{y}+\alpha)$
@@
Solving the maximin game.
We solve the relaxed adversarial game given in Eq. 10 by alternate gradient-based optimization (Goodfellow et al., 2016). Concretely, we alternate between updating the predictor’s parameters ${\boldsymbol{\theta}}$ and the projection matrix ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}$. Updates to ${\boldsymbol{\theta}}$ are performed with gradient descent, and updates to ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{P}}$ are performed with gradient ascent, including a projection onto the Fantope to ensure that the constraint is met. For the Fantope projection step, we use Vu et al.’s (2013) algorithm, the details of which are restated in Ravfogel et al. (2022).
@@ 5 Effect on Concept Encoding
5.1 Pre-image Recovery: Same Adversary
5.2 Pre-image Recovery: Kernel Transfer
@@ Results.
In Table 3, we report the test statistic and the $p$-value for the third test using the names of scientific and artistic fields to represent gender-biased words. [^10] See § A.6 for results of the other two biased word association tests. For all kernels, we observe a significant drop in the test statistic, from the original value of 1.56 to around 0.74. This suggests that the intervention significantly decreases the association between female and male names and stereotypically biased words. Notably, the reduction is similar for all kernels, including the linear one. While non-linear erasure is more effective in neutralizing gender against adversarial recovery, linear and non-linear methods perform equally well according to this bias association test. This finding highlights the importance of measuring different manifestations of a concept when using concept neutralization as a bias mitigation method.
@@ Appendix A Related Work
A.1 A Representer Lemma for Kernelized Maximin Games
@@
$\max_{\boldsymbol{w},\in\mathcal{H}}\min_{{\boldsymbol{\theta}}\in\mathcal{H}}\sum_{n=1}^{N}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\ell}\left(y_{n},\,\Big{\langle}{\boldsymbol{\theta}},\left(\mathrm{I}-\frac{{\boldsymbol{w}\boldsymbol{w}^{\top}}}{\boldsymbol{w}^{\top}\boldsymbol{w}}\right)\boldsymbol{\Phi}(\boldsymbol{x}_{n})\Big{\rangle}\right).$ (12)
@@
$=\boldsymbol{\Phi}(\mathbf{x}_{n})-\frac{\boldsymbol{w}_{U}\boldsymbol{w}_{U}^{\top}}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}\boldsymbol{\Phi}(\mathbf{x}_{n})-\underbrace{\frac{\boldsymbol{w}_{\perp U}\boldsymbol{w}_{\perp U}^{\top}}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}}_{=0}\boldsymbol{\Phi}(\mathbf{x}_{n})$ (13d)
$=\boldsymbol{\Phi}(\mathbf{x}_{n})-\boldsymbol{w}_{U}\frac{\boldsymbol{w}_{U}^{\top}\boldsymbol{\Phi}(\mathbf{x}_{n})}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}.$ (13e)
@@
$\Big{\langle}{\boldsymbol{\theta}},\left(\mathrm{I}-{\boldsymbol{w}\boldsymbol{w}^{\top}}\right)\boldsymbol{\Phi}(\boldsymbol{x}_{n})\Big{\rangle}={\boldsymbol{\theta}}_{U}^{\top}\boldsymbol{\Phi}(\mathbf{x}_{n})-{\boldsymbol{\theta}}_{U}^{\top}\boldsymbol{w}_{U}\frac{\boldsymbol{w}_{U}^{\top}\boldsymbol{\Phi}(\mathbf{x}_{n})}{\boldsymbol{w}^{\top}{\boldsymbol{w}}}.$ (14)
author=ryan,color=violet!40,size=,fancyline,caption=,]We are missing a denominator in 14. Thus, for any ${\boldsymbol{\theta}},\boldsymbol{w},\in\mathcal{H}$ there exists a ${\boldsymbol{\theta}}_{U},\boldsymbol{w}_{U}\in U$ that yields the same value of the objective as ${\boldsymbol{\theta}},\boldsymbol{w}$. Now, we can parameterize ${\boldsymbol{\theta}}_{U}$ and $\boldsymbol{w}_{U}$ as
@@
${\boldsymbol{\theta}}_{U}$ $=\sum_{n=1}^{N}\beta_{n}\boldsymbol{\Phi}(\mathbf{x}_{n}).$ (16)
for real coefficients $\boldsymbol{\alpha}\in\mathbb{R}^{N}$ and $\boldsymbol{\beta}\in\mathbb{R}^{N}$. We conclude that for any local optimum ${\boldsymbol{\theta}}^{\star},\boldsymbol{w}^{\star}\in\mathcal{H}$, the projection of ${\boldsymbol{\theta}}^{\star}$ and $\boldsymbol{w}^{\star}$ onto $U$ yields a local optimum with the same value. ∎
A.2 Kernelization of the Maximin Game
We show that the game Eq. 1 can be kernelized for the case $k=1$, i.e., a setting where the matrix $\mathcal{P}_{k}$ removes a one-dimensional subspace. Specifically, we will show that the product $\langle{\boldsymbol{\theta}},\mathrm{P}\boldsymbol{\Phi}(\boldsymbol{x}_{n})\rangle$ in Eq. 1 can be expressed as a function of the kernel ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\cdot,\cdot)$.